OpenStudy (anonymous):

7x - 4y = 33 6x - 7y = 14

OpenStudy (anonymous):

the answer is... idk the answer. haha :)

OpenStudy (anonymous):

its quite easy x=33+4y/7 insert value of x into 6x - 7y = 14 it will yield value of y

OpenStudy (anonymous):

repeat the process for x

OpenStudy (anonymous):

7x-4y=33....A 6x-7y=14...........B use equation A to find out one variable and then put its value in equation B......get it

OpenStudy (anonymous):

like 7x-4y=33 7x=33+4y x=33+4y/7 put this value of x in equation B after finding y put its value in equation A you will find x @dezzy hope you get it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

solve equation 1 for x 7x=33+4y x=(33+4y)/7 put in second equation 6((33+4y)/7)-7y=14 (198+24y)/7-7y=14 198+24y-49y=98 198-25y=98 25y=198-94 y=100/25=4 y=4 since x=(33+4y)/7 so use y=4 x=(33+4(4)/7 x=49/7 x=7 x=7 y=4

OpenStudy (anonymous):

thanks sami

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

you r welcum